the path difference between two interfeering waves ata point on the screen is x/3, where x is wave leength , from the central maximum, the ratio of intensity at this point and at the central fringe will be

  • the path difference between two interfeering waves ata point on the screen is x/3, where x is wave leength , from the central maximum, the ratio of intensity at this point and at the central fringe will be

Grade:12th pass

1 Answers

Neeti
8 years ago
\Deltax/\lambda = \Delta\Phi/2\pi so we have path difference delta x = x/3 where  x= \lambda = wavelength. 
From here we get phase difference between the two =\Delta \Phi
Now when phase difference is \Phi the resultant intensity can be calculated by :
Inet =I1+ I2 + sq.root (I1xI2)cos \Phi (here we assume I1 = I2 = I ) and intensity at central maxima = 4I (since \Delta\Phi = zero for center, Inet = 4I ) 
 
So you get the ratio. ( Inet / 4I )